# Improved Moving-Puncture Techniques for Compact Binary Simulations

### v2

@@ Abstract
To fully unlock the scientific potential of upcoming gravitational wave (GW) interferometers, numerical relativity (NR) simulation accuracy will need to be greatly enhanced. We present three infrastructure-agnostic improvements to the moving-puncture approach for binary black hole (BBH) simulations, aimed at greatly reducing constraint violation and improving GW predictions. Although these improvements were developed within the highly efficient NR code BlackHoles@Home, we demonstrate their effectiveness in the widely-adopted Einstein Toolkit/Carpet AMR framework. Our improvements include a modified Kreiss-Oliger dissipation prescription, a Hamiltonian-constraint-damping adjustment to the BSSN equations, and an extra term to the 1+log lapse evolution equation that slows the development of the sharp lapse feature, which dominates numerical errors in BBH simulations. With minimal increase in computational cost, these improvements greatly reduce GW noise, enabling the extraction of high-order GW modes previously obscured by numerical noise. They also improve convergence properties, reduce Hamiltonian (momentum) constraint violations in the strong-field region by roughly two (three) orders of magnitude, and in the GW-extraction zone by five (two) orders of magnitude. To promote community adoption, we have open-sourced the improved Einstein Toolkit thorn BaikalVacuum used in this work. Although our focus is on BBH evolutions and the BSSN formulation, these improvements may also benefit compact binary simulations involving matter and other formulations, a focus for future investigations.
@@ I Introduction
When it comes to solving Einstein’s equations in NR, the moving-puncture approach is the most popular method. This approach generally combines the BSSN Baumgarte and Shapiro (1998); Shibata and Nakamura (1995), CCZ4 Alic et al. (2012) or Z4c Bernuzzi and Hilditch (2010) 3+1 decompositions of Einstein’s equations with the moving-puncture gauge conditions Baumgarte and Shapiro (1998); Shibata and Nakamura (1995); Nakamura et al. (1987); Alcubierre et al. (2003), enabling the evolution of black holes (BHs) that traverse the numerical grid without the need to excise their interiors Baker et al. (2006); Campanelli et al. (2006). However, mixing the moving-puncture technique with Cartesian AMR grids results in significant numerical noise that negatively impacts the accuracy of compact binary simulations.
@@
This paper introduces improvements to the moving-puncture approach, aimed at curtailing the development of sharp features in the lapse and mitigating their impact.
@@
Although our focus is primarily on the impact of these improvements on BBH evolutions with the BSSN formalism, we expect that they will also be beneficial in other compact binary evolutions and formalisms that use the MP technique. Determining whether our improved techniques significantly improve the fidelity of binary neutron star (BNS) simulations will be a focus of an upcoming paper. In addition, formulations like Z4c have proven quite effective in reducing constraint violations in BNS simulations Hilditch et al. (2013). Comparing BNS evolutions performed with our improved BSSN moving-puncture approach to Z4 techniques widely used across the community will also be a focus of future work.
@@ II.1 BSSN Evolution Equations
where $\phi$ represents the conformal factor. Instead of evolving $\phi$, we follow previous works (e.g., Tichy and Marronetti (2007); Marronetti et al. (2008)) and evolve $W=e^{-2\phi}$, as $W$ exhibits greater smoothness near puncture BHs, resulting in reduced truncation errors. To this end, derivatives of the conformal factor $\phi$ appearing in the BSSN evolution equations are calculated with respect to the evolved variable $W$ via $\partial_{\mu}W=-2e^{-2\phi}\partial_{\mu}\phi\Rightarrow\partial_{\mu}\phi=-\partial_{\mu}W/(2W)$, and so forth for second derivatives.
@@
Building on these definitions, we adopt the following time-evolution equations:
@@
$+W^{2}\left\{-2\alpha\bar{D}_{i}\bar{D}_{j}\phi+4\alpha\bar{D}_{i}\phi\bar{D}_{j}\phi\right.$
@@ II.2 Moving-Puncture Gauge Evolution Equations
To close the system of time-evolution equations, we choose the moving-puncture gauge conditions, which include the 1+log slicing condition for the lapse and the $\Gamma$-driving (here, $\Lambda$ takes the place of $\Gamma$) condition for the shift:
@@ V.2 Strong-Field Region: Effect of Improvements on Constraint Violations
where $A$ is related to the $n+1$st derivative of the term in $C_{i}$ that dominates the truncation error. So long as $C_{i}$ is known at at least two resolutions, both $n$ and $A$ can be computed.
@@ VI Conclusions and Future Work
When combining all three improvements (All3), performance is significantly better than any single enhancement, cutting down Hamiltonian (momentum) constraint violations in the strong-field area by roughly two (three) orders of magnitude and in the GW-extraction zone by five (two) orders of magnitude. Further, the convergence of constraint violations improves, suggesting that these improvements become even more pronounced at higher resolutions.
@@ VII Acknowledgments
We would like to thank S. Cupp, P. Diener, R. Haas, B. J. Kelly, and L. R. Werneck for their comments on early drafts of this manuscript. We would also like to thank J. G. Baker, B. Brügmann, W. East, E. Hirschmann, C. O. Lousto, D. Neilsen, V. Paschalidis, and Y. Zlochower for helpful discussions in the preparation of this manuscript. Z.B.E. gratefully acknowledges support from NSF awards PHY-2110352, OAC-2227105, AST-2227080, as well as NASA awards ISFM-80NSSC18K0538 and TCAN-80NSSC18K1488. This research made use of the resources of the High Performance Computing Center at Idaho National Laboratory, which is supported by the Office of Nuclear Energy of the U.S. Department of Energy and the Nuclear Science User Facilities under Contract No. DE-AC07-05ID14517. In addition, it made use of the Falcon Idaho C3+3 Collaboration (2022) supercomputer, operated by the Idaho C3+3 Collaboration.

### v3

@@ Abstract
To fully unlock the scientific potential of upcoming gravitational wave (GW) interferometers, numerical relativity (NR) simulation accuracy will need to be greatly enhanced. We present three infrastructure-agnostic improvements to the moving-puncture approach for binary black hole (BBH) simulations, aimed at greatly reducing constraint violation and improving GW predictions. Although these improvements were developed within the highly efficient NR code BlackHoles@Home, we demonstrate their effectiveness in the widely-adopted Einstein Toolkit/Carpet AMR framework. Our improvements include a modified Kreiss-Oliger dissipation prescription, a Hamiltonian-constraint-damping adjustment to the BSSN equations, and an extra term to the 1+log lapse evolution equation that slows the development of the sharp lapse feature, which dominates numerical errors in BBH simulations. With minimal increase in computational cost, these improvements greatly reduce GW noise, enabling the extraction of high-order GW modes previously obscured by numerical noise. They also improve convergence properties near and inside the convergent regime, reduce Hamiltonian (momentum) constraint violations in the strong-field region by roughly two (three) orders of magnitude, and in the GW-extraction zone by five (two) orders of magnitude. To promote community adoption, we have open-sourced the improved Einstein Toolkit thorn BaikalVacuum used in this work. Although our focus is on BBH evolutions and the BSSN formulation, these improvements may also benefit compact binary simulations involving matter and other formulations, a focus for future investigations.
@@ I Introduction
When it comes to solving Einstein’s equations in NR, the moving-puncture approach Campanelli et al. (2006); Baker et al. (2006) is the most popular method. This approach generally combines the BSSN Baumgarte and Shapiro (1998); Shibata and Nakamura (1995), CCZ4 Alic et al. (2012) or Z4c Bernuzzi and Hilditch (2010) 3+1 decompositions of Einstein’s equations with the moving-puncture gauge conditions Alcubierre et al. (2003); Campanelli et al. (2006); Baker et al. (2006), enabling the evolution of black holes (BHs) that traverse the numerical grid without the need to excise their interiors. However, mixing the moving-puncture technique with Cartesian AMR grids results in significant numerical noise that negatively impacts the accuracy of compact binary simulations.
@@
We emphasize that this discussion has been limited to Cartesian AMR grids not set to refine at the sharp gauge features, as this approach is widely used in the field. Improved refinement algorithms or grid structures may result in significantly less constraint violation and GW noise than what is reported here.
This paper introduces improvements to the moving-puncture approach, aimed at curtailing the development of sharp features in the lapse and mitigating their impact. Our first two improvements focus on decreasing the strength of low-pass filters near punctures—where sharp fields are desirable—and increasing the strength far away.
@@
Although our focus is primarily on the impact of these improvements on BBH evolutions with the BSSN formalism, we expect that they will also be beneficial in other compact binary evolutions and formalisms that use the moving-puncture technique. Determining whether our improved techniques significantly improve the fidelity of binary neutron star (BNS) simulations will be a focus of an upcoming paper. In addition, formulations like Z4c have proven quite effective in reducing constraint violations in BNS simulations Hilditch et al. (2013). Comparing BNS evolutions performed with our improved BSSN moving-puncture approach to Z4 techniques widely used across the community will also be a focus of future work.
@@ II.1 BSSN Evolution Equations
where $\phi$ represents the conformal factor. Instead of evolving $\phi$, we follow previous works (e.g., Tichy and Marronetti (2007); Marronetti et al. (2008)) and evolve $W=e^{-2\phi}$, as $W$ exhibits greater smoothness near puncture BHs, resulting in reduced truncation errors (also see discussion in Campanelli et al. (2006)). To this end, derivatives of the conformal factor $\phi$ appearing in the BSSN evolution equations are calculated with respect to the evolved variable $W$ via $\partial_{\mu}W=-2e^{-2\phi}\partial_{\mu}\phi\Rightarrow\partial_{\mu}\phi=-\partial_{\mu}W/(2W)$, and so forth for second derivatives.
@@
Building on these definitions, we adopt the following time-evolution equations, replacing all occurrences of $\phi$ and its derivatives with equivalent expressions in terms of $W$:
@@
$+e^{-4\phi}\left\{-2\alpha\bar{D}_{i}\bar{D}_{j}\phi+4\alpha\bar{D}_{i}\phi\bar{D}_{j}\phi\right.$
@@ II.2 Moving-Puncture Gauge Evolution Equations
To close the system of time-evolution equations, we choose the moving-puncture gauge conditions Alcubierre et al. (2003); Campanelli et al. (2006); Baker et al. (2006), which include the 1+log slicing condition for the lapse and the $\Gamma$-driving (here, $\Lambda$ takes the place of $\Gamma$) condition for the shift:
@@
Variations in moving-puncture gauge conditions, such as a first-order implementation of the shift condition, do exist. Although we did not explore these alternatives in our study, we direct the reader to van Meter et alvan Meter et al. (2006) for a comprehensive analysis of different implementations. Our choice of gauge conditions was primarily motivated by their finding that incorporating advection terms into the lapse and shift conditions leads to well-behaved characteristic speeds.
@@ V.2 Strong-Field Region: Effect of Improvements on Constraint Violations
where $A$ is related to the $(n+1)$st derivative of the term in $C_{i}$ that dominates the truncation error. So long as $C_{i}$ is known at at least two resolutions, both $n$ and $A$ can be computed.
@@ VI Conclusions and Future Work
When combining all three improvements (All3), performance is significantly better than any single enhancement, cutting down Hamiltonian (momentum) constraint violations in the strong-field area by roughly two (three) orders of magnitude and in the GW-extraction zone by five (two) orders of magnitude. Further, the convergence of constraint violations improves near and inside the convergent regime, suggesting that these improvements become even more pronounced at higher resolutions.
@@ VII Acknowledgments
We would like to thank S. Cupp, P. Diener, R. Haas, B. J. Kelly, and L. R. Werneck for their comments on early drafts of this manuscript. We would also like to thank J. G. Baker, W. K. Black, B. Brügmann, W. East, E. Hirschmann, C. O. Lousto, D. Neilsen, V. Paschalidis, and Y. Zlochower for helpful discussions in the preparation of this manuscript. Z.B.E. gratefully acknowledges support from NSF awards PHY-2110352, OAC-2227105, AST-2227080, as well as NASA awards ISFM-80NSSC18K0538 and TCAN-80NSSC18K1488. This research made use of the resources of the High Performance Computing Center at Idaho National Laboratory, which is supported by the Office of Nuclear Energy of the U.S. Department of Energy and the Nuclear Science User Facilities under Contract No. DE-AC07-05ID14517. In addition, it made use of the Falcon Idaho C3+3 Collaboration (2022) supercomputer, operated by the Idaho C3+3 Collaboration.